# TraCS: Trajectory Collection in Continuous Space under Local Differential Privacy

### v2

@@ 3.5.2 Impact of the Size of $\mathcal{S}$
3.5.3 Comparison with Geo-indistinguishability
Geo-indistinguishability (Geo-Ind for brevity, also called metric privacy in recent works) [2] is a more general (or weaker) version of LDP. From the perspective of Geo-ind, LDP force the proximity-aware function $d(x,y)$ in Geo-ind to be $d(x,y)=1$ for all $x,y\in S$, ignoring that different $x$ and $y$ have different $d(x,y)$ in the context of their locations. Conversely, if the upper bound $\max_{x,y\in S}d(x,y)\leq d^{*}$ holds in $S$, then TraCS also satisfies $d^{*}\varepsilon$-Geo-ind. However, this trivial satisfaction of $d^{*}\varepsilon$-Geo-ind is highly suboptimal, since it uses the global upper bound of $d(x,y)$ rather than the fine-grained, actual values of $d(x,y)$ for different location pairs of $x$ and $y$.
@@ 3.5.3 Comparison with Geo-indistinguishability
3.5.4 Other LDP Mechanisms for Bounded and Continuous Space
@@
In this paper, we leverage piecewise-based mechanisms for bounded numerical domains to design the direction and distance perturbations in TraCS. Other LDP mechanisms for bounded numerical domains can also be incorporated into TraCS, but most of them may require redesign to suit the direction perturbation, as the circular domain $\mathcal{D}_{\varphi}=[0,2\pi)$ has a different distance metric than linear domains. One exception is the Purkayastha mechanism [33], which is designed for sphere $\mathbb{S}^{n-1}$, e.g. circular lines when $n=2$. It can be an alternative to $\mathcal{M}_{\circ}$, but not be directly applicable to $S\subset\mathbb{R}^{2}$. We don’t leverage it for direction perturbation because of its sampling complexity (Section 3.4.2 in [33]) and higher MSE.
@@
Another category of LDP mechanisms for bounded numerical domains is truncated mechanisms, including truncated Laplace [22, 15] and truncated Gaussian mechanisms [4]. These mechanisms modify Laplace and Gaussian distributions by redesigning distributions on a bounded domain, e.g. $[0,1)$. However, truncated mechanisms cannot preserve the same privacy level as their original counterparts [22, 15, 4]. Determining their new privacy level is non-trivial and typically relies on numerical testing algorithms [15, 4] rather than closed-form solutions. Compared with truncated mechanisms, piecewise-based mechanisms are explicitly designed for LDP in bounded numerical domains, making them more effective.
3.5.5 More Extensions
Appendix C.4 provides two more extensions of TraCS and experimental results, including (i) other piecewise-based mechanisms and (ii) other shapes of location space. The experimental evaluations are included in Appendix C.6.
4 Evaluation
@@
4.1 Results on Continuous Space
The strawman method, our extension of $k$-$R$ for direction perturbation, can be combined with the distance perturbation mechanism in TraCS-D to ensure LDP in continuous spaces $\mathcal{S}\subset\mathbb{R}^{2}$. For the name simplicity, we still refer to this extended method as the strawman method and compare it with TraCS in continuous space.
4.1.1 Setup
TraCS requires the following parameters: the location space $\mathcal{S}$, the sensitive trajectory $\mathcal{T}$, the privacy parameter $\varepsilon$, and $\varepsilon_{d}$ for direction perturbation in TraCS-D. The strawman method also requires an instance of $k$ in $k$-$R$. We use configurations:
  • Location space $\mathcal{S}$: We consider synthetic and real-world settings. (i) The default location space is $0,1)\times[0,1)$. Without loss of generality, we consider a larger space $\mathcal{S}=[0,2)\times[0,10)$ for comparison. We generate $100$ random trajectories for each location space as synthetic datasets. This setup excludes dataset-specific effects to emphasize algorithmic performance. (ii) Areas of TKY and CHI: These two trajectory datasets were collected in Tokyo and Chicago, extracted from the Foursquare dataset [34] and Gowalla dataset [7], respectively. Locations in TKY are within area $139.47,139.90)\times[35.51,35.86)$ and CHI within area $87.9,-87.5)\times[41.6,42.0)$. We use the first $100$ trajectories in each dataset.
  • Privacy parameter $\varepsilon_{d}$: We set $\varepsilon_{d}=\varepsilon\pi/(\pi+1)$ to heuristically balance the perturbation in $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$ according to their domain sizes.
@@ 4.1.1 Setup
Trajectory utility metrics. We follow the Euclidean distance to measure the utility of the perturbed trajectory [37]. The distance between two locations $\tau_{i}$ and $\tau^{\prime}_{i}$ is defined as $\|\tau_{i}-\tau^{\prime}_{i}\|_{2}$. Therefore, given a sensitive trajectory $\mathcal{T}$ and a perturbed trajectory $\mathcal{T}^{\prime}$, the average error (AE) among all locations in the trajectory is:
@@
where $|\mathcal{T}|$ is the number of locations in the trajectory. A smaller $AE$ indicates better utility of the perturbed trajectory. Many other trajectory utility metrics, such as Preservation Range Query [9, 37], i.e. whether each perturbed location is within a $\delta$ range of the corresponding sensitive location, are applications of $AE$ and can be derived from it. We defer these results to Appendix C.7 for brevity.
4.1.2 Error Comparison on Synthetic Datasets
@@
Figure 4: Comparison on synthetic datasets.
Figure 5: Condition TraCS-D outperforms TraCS-C.
Figure 6: TraCS-D vs $k$-$R$ + uniform sampling.
Figure 4 presents the error comparison between TraCS and the strawman method. We can observe the error advantage of TraCS over the strawman method as the privacy parameter $\varepsilon$ increases. Specifically, in Figure 4(a), TraCS-C exhibits smaller $AE$ when $\varepsilon\gtrsim 3$, and TraCS-D exhibits smaller $AE$ when $\varepsilon\gtrsim 5$. When the location space expands to $0,2)\times[0,10)$, i.e. Figure 4(b), the advantage of TraCS becomes more significant. Among TraCS, TraCS-C consistently has a smaller $AE$ than TraCS-D. This is because TraCS-D often deals with larger distance spaces $\mathcal{D}_{r(\varphi)}$, which also magnifies the error of direction perturbation. Statistically, the mean $AE$ of TraCS-D is $91.1\%$ of the strawman method in Figure 4(a) and $86.6\%$ in Figure 4(b) across all the $\varepsilon$ values. For TraCS-C the corresponding ratios are $75.5\%$ in Figure 4(a) and $64.0\%$ in Figure 4(b).
@@
Figure 7: Comparison on real-world datasets.
4.1.3 Error Comparison on Real-world Datasets
Figure 7 shows the error comparison on the TKY and CHI areas. The trends mirror those observed on the synthetic data. Across all $\varepsilon$ values, the mean $AE$ of TraCS-D is $96.8\%$ of the strawman method on TKY and $94.5\%$ on CHI; for TraCS-C the corresponding ratios are $89.7\%$ and $61.2\%$. The performance gap is smaller on TKY than on CHI, because TKY’s location space is more compact, which lowers the error values for all perturbation methods.
4.1.4 Choose TraCS-D or TraCS-C
@@
4.2 Results on Discrete Space
TraCS can be applied to discrete spaces by rounding the perturbed locations to the nearest discrete points. Notably, the privacy guarantee remains valid for the continuous space and hence is also valid for the discrete space, independent of the discretization. We will evaluate the performance of TraCS on discrete spaces and compare it with existing methods.
4.2.1 Setup
@@
  • NGram [9]: It is often impractical to acquire reachability knowledge in practice [37]. We instead consider a strong reachability constraint, i.e. from each location, the next location can only be within a distance of $0.75\cdot\hat{a}$, [^4] While the original NGram employs much stricter time-reachability constraints (e.g. the next location must be within 0.10.1h $\cdot$ 44km/h), this constraint is also known to the adversary, which undermines the privacy guarantee. TraCS are LDP mechanisms guaranteeing privacy for the entire domain and do not incorporate such constraints. where $\hat{a}$ is the maximal length of the location space. This constraint significantly reduces the location space for perturbation, especially for the locations near the boundary.
  • L-SRR [29]: We use group number $g=3$, the same as the empirical optimal value in their paper and code. Locations in the first group are within $0.3\cdot\hat{a}$ distance from the sensitive location, where $\hat{a}$ is the maximal length of the location space. The second group is within $0.6\cdot\hat{a}$ distance from the sensitive location, and the remaining locations are in the third group.
  • ATP [37]: We set the $k$-$R$ parameter $k=6$ in ATP. The privacy parameter $\varepsilon$ in original ATP is for the entire trajectory, we distribute it to each location’s perturbation according to their algorithms.
@@
  • Synthetic datasets: We uniformly discretize $0,1)\times[0,1)$ into $m$ discrete points and treat the $m$ points as the discrete location space. Specifically, we set $m=10\times 10$ and $m=60\times 60$ for comparison. Along with them, we generate $100$ random trajectories with $100$ locations each as synthetic datasets.
  • TKY and CHI: We treat every location in the TKY and CHI datasets as a discrete point, which together constitute the discrete location space. TKY contains $7,798$ discrete locations; CHI contains $1,000$ discrete locations. We use the first $100$ trajectories for evaluation.
For the TKY and CHI datasets, we perform the perturbation $5$ times for each trajectory and compute the average $AE$, which is the same setup as in ATP [37].
4.2.2 Error Comparison on Synthetic Datasets
@@
Figure 8: Comparison on synthetic datasets. $m$ is the number of discrete points in the location space.
@@
4.2.3 Error Comparison on Real-world Datasets
@@
Reasons for the results. (i) ATP performs exceptionally well when $\varepsilon$ is small, particularly on the TKY dataset. This is because this dataset contains many “condensed” locations, i.e. no other locations near the trajectory. In such cases, if the bi-direction perturbation of ATP is accurate, the perturbed location is error-free. (ii) The $AE$ of ATP does not decrease significantly as $\varepsilon$ increases, which is consistent with their paper. NGram shows a similar trend, especially on the TKY dataset. This is due to the low performance of the Exponential mechanism for large location spaces, particularly when the locations are concentrated in a small area. In such cases, the score function struggles to differentiate the scores among the locations. L-SRR exhibits a similar trend. (iii) TraCS exhibits larger $AE$ than them when $\varepsilon$ is small because it focuses on the rectangular area formed by the longitude and latitude, instead of a set of discrete locations in the cities. Therefore, the perturbed locations may be far from the city center (where locations are condensed and benefits discrete mechanisms) when $\varepsilon$ is small.
@@
Figure 9: Comparison on real-world datasets. TraCS operates on rectangular areas encompassing the city, which may result in larger AEAE when $\varepsilon$ is small.
Table 2: Time cost comparison (in milliseconds).
ATP NGram L-SRR TraCS-D TraCS-C
Total $145.7$ $100.9$ $6.2$ $0.06$ $0.05$
Perturb $125.8$ $92.8$ $0.003$ $0.018$ $0.003$
4.2.4 Time Cost Comparison

### v3

@@ 3.5.4. Impact of the Size of $\mathcal{S}$
3.5.5. Comparison with Geo-indistinguishability
Geo-indistinguishability (DBLP:conf/ccs/AndresBCP13) (Geo-Ind for brevity, also called metric privacy (DBLP:journals/popets/QiuLPSX25) in recent works) is a more general (or weaker) version of LDP. From the perspective of Geo-ind, LDP force the proximity-aware function $d(x,y)$ in Geo-ind to be $d(x,y)=1$ for all $x,y\in S$, ignoring that different $x$ and $y$ have different $d(x,y)$ in the context of their locations. Conversely, if the upper bound $\max_{x,y\in S}d(x,y)\leq d^{*}$ holds in $S$, then TraCS also satisfies $d^{*}\varepsilon$-Geo-ind. However, this trivial satisfaction of $d^{*}\varepsilon$-Geo-ind is highly suboptimal, since it uses the global upper bound of $d(x,y)$ rather than the fine-grained, actual values of $d(x,y)$ for different location pairs of $x$ and $y$.
@@ 3.5.5. Comparison with Geo-indistinguishability
3.5.6. Other LDP Mechanisms for Bounded and Continuous Space
@@
In this paper, we employ utility-optimized piecewise-based mechanisms (OGPM) (DBLP:journals/popets/ZhengMH25) for bounded numerical domains to design the direction and distance perturbations in TraCS. Other LDP mechanisms for bounded numerical domains can also be incorporated into TraCS, but most of them may require redesign to suit the direction perturbation, as the circular domain $\mathcal{D}_{\varphi}=[0,2\pi)$ has a different distance metric than linear domains. One exception is the Purkayastha mechanism (DBLP:conf/ccs/WeggenmannK21), which is designed for sphere $\mathbb{S}^{n-1}$, e.g. circular lines when $n=2$. It can be an alternative to $\mathcal{M}_{\circ}$, but not be directly applicable to $S\subset\mathbb{R}^{2}$. We don’t leverage it for direction perturbation because of its sampling complexity (Section 3.4.2 in (DBLP:conf/ccs/WeggenmannK21)) and higher MSE (DBLP:journals/popets/ZhengMH25).
@@
Another category of LDP mechanisms for bounded numerical domains is truncated mechanisms, including truncated Laplace (DBLP:journals/tdp/000619; DBLP:journals/jpc/HolohanABA20) and truncated Gaussian mechanisms (DBLP:journals/jpc/ChenH24). These mechanisms modify Laplace and Gaussian distributions by redesigning distributions on a bounded domain, e.g. $[0,1)$. However, truncated mechanisms cannot preserve the same privacy level as their original counterparts (DBLP:journals/tdp/000619; DBLP:journals/jpc/HolohanABA20; DBLP:journals/jpc/ChenH24). Determining their new privacy level is non-trivial and typically relies on numerical testing algorithms (DBLP:journals/jpc/HolohanABA20; DBLP:journals/jpc/ChenH24) rather than closed-form solutions. Compared with truncated mechanisms, piecewise-based mechanisms are explicitly designed for LDP in bounded numerical domains, making them more effective and easier to analyze.
3.5.7. Other Piecewise-based Mechanisms
In Definition 1 and Definition 3, we use specific instantiations for the parameters $p$ and $[l,r)$ in the piecewise-based mechanism. Besides these instantiations, TraCS can be integrated with any other piecewise-based mechanisms (DBLP:conf/icde/WangXYZHSS019; DBLP:conf/sigmod/Li0LLS20; DBLP:journals/tmc/MaZW24; 10735696; DBLP:journals/iotj/ZhaoZYWWLNL21) by redesigning them. For example, SW (DBLP:conf/sigmod/Li0LLS20; DBLP:journals/popets/ZhengMH25) is a famous mechanism proposed for distribution estimation; it is defined on $0,1)\to[-b,1+b)$, with $b$ as a parameter. It uses different $p$ and $[l,r)$, and employs a strategy of maximizing the mutual information between the perturbed and sensitive data to determine $b$. SW can be redesigned (DBLP:journals/popets/ZhengMH25) for the direction perturbation and distance perturbation in TraCS-D and TraCS-C. Appendix C.6 provides the detailed redesign of SW for them and shows intuitive comparisons with Definition 1 and Definition 3, and Appendix C.7 provides a comparison.
3.5.8. Other Shapes of Location Space
The design of LDP mechanisms is space-specific, implying that extensions to other shapes of location space are not straightforward.
By post-processing. A practical workaround for non-rectangular continuous domains is to post-process perturbed locations so they lie inside the target region. If a perturbed location falls outside the shape, project it to the nearest point inside the shape (for polygons this is the closest point on the boundary). Because the shape is public and projection is a data-independent post-processing step, it does not weaken the LDP guarantee. This approach is especially useful for irregular shapes or unions of multiple cities.
Direct extension. Directly extending TraCS to other shapes requires decomposing the domain into independent subspaces that are compatible with our perturbation primitives. For TraCS, the challenge lies in determining the size of each subspace, i.e. $|\mathcal{D}_{\varphi}|$, $|\mathcal{D}_{r(\varphi)}|$, $|\mathcal{D}_{a}|$, and $|\mathcal{D}_{b}|$. Among them, $|\mathcal{D}_{r(\varphi)}|$ is not apparent but can be calculated for rectangular location spaces. Although calculating $|\mathcal{D}_{r(\varphi)}|$ is generally non-trivial for irregular shapes, TraCS can be extended to parallelogram location spaces by a linear transformation to a rectangular space. Meanwhile, TraCS-D can be extended to circular regions, which are widely used in relative-location representations (e.g. radar and sonar systems). Extending TraCS-C to circular regions is more challenging, since such regions do not admit a natural Cartesian decomposition into independent coordinates. Appendix C.7 reports experimental results of TraCS-D on circular location spaces.
4. Evaluation
@@
4.1. Results on Continuous Space
The strawman method—our extension of $k$-$R$ for direction perturbation—can be combined with the distance perturbation mechanism in TraCS-D to ensure LDP in continuous spaces $\mathcal{S}\subset\mathbb{R}^{2}$. For simplicity, we continue to refer to this extended approach as the strawman method, and compare it with TraCS in continuous space. We also include the planar Laplace mechanism (DBLP:conf/ccs/AndresBCP13) ++ clipping (truncation) as a baseline. This baseline perturbs each location by adding noise drawn from a 2D planar Laplace distribution, and then truncates the output to the location space $\mathcal{S}$ if it falls outside. [^7] Appendix C.8 provides more details on the planar Laplace mechanism + clipping.
4.1.1. Setup
TraCS requires the following parameters: the location space $\mathcal{S}$, the sensitive trajectory $\mathcal{T}$, the privacy parameter $\varepsilon$, and $\varepsilon_{d}$ for direction perturbation in TraCS-D. The strawman method also requires an instance of $k$ in $k$-$R$. We use configurations as follows for the experiments.
  • Location space $\mathcal{S}$: We consider synthetic and real-world settings. (i) The default location space is $0,1)\times[0,1)$. Without loss of generality, we consider a larger space $\mathcal{S}=[0,2)\times[0,10)$ for comparison. We generate $100$ random trajectories for each location space as synthetic datasets. Specifically, each trajectory is generated by randomly sampling a sequence of $100$ points in $\mathcal{S}$. This setup excludes dataset-specific effects to emphasize algorithmic performance. (ii) Areas of TKY and CHI: These two trajectory datasets were collected in Tokyo and Chicago, extracted from the Foursquare dataset (DBLP:journals/tsmc/YangZZY15) and Gowalla dataset (DBLP:conf/kdd/ChoML11), respectively. Locations in TKY are within area $139.47,139.90)\times[35.51,35.86)$ and CHI within area $87.9,-87.5)\times[41.6,42.0)$. These two areas are visually shown in Figure 18. We use the first $100$ trajectories in each dataset.
  • Privacy parameter $\varepsilon_{d}$: We set $\varepsilon_{d}=\varepsilon\pi/(\pi+1)$ to heuristically balance the perturbation in $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$ according to their domain sizes. [^8] Appendix C.9 provides more discussion on setting $\varepsilon_{d}$.
@@ 4.1.1. Setup
Trajectory utility metrics. We follow the Euclidean distance to measure the utility of the perturbed trajectory (DBLP:journals/pvldb/Zhang [^0] H23; DBLP:journals/pvldb/CunninghamCFS21; DBLP:conf/ccs/WangH0QH22). The distance between two locations $\tau_{i}$ and $\tau^{\prime}_{i}$ is defined as $\|\tau_{i}-\tau^{\prime}_{i}\|_{2}$. Therefore, given a sensitive trajectory $\mathcal{T}$ and a perturbed trajectory $\mathcal{T}^{\prime}$, the average error (AE) among all locations in the trajectory is:
@@
where $|\mathcal{T}|$ is the number of locations in the trajectory. A smaller AE indicates better trajectory utility. We compute the average AE across all trajectories in the dataset for comparison. Although AE is not a perfect utility measure, e.g. it may not fully reflect the requirements of specific downstream tasks, it remains the commonly used metric for evaluating the utility of perturbed trajectories (DBLP:journals/popets/MirandaPascualGPFS23; DBLP:journals/eswa/ZhaoPC20; DBLP:journals/pvldb/Zhang [^0] H23; DBLP:journals/pvldb/CunninghamCFS21; DBLP:conf/ccs/WangH0QH22). Moreover, many other trajectory utility metrics are directly or indirectly related to AE, such as the preservation of range queries and hotspots (DBLP:journals/pvldb/CunninghamCFS21; DBLP:journals/pvldb/Zhang [^0] H23). For brevity, we defer these additional metrics to Section 4.2.5.
4.1.2. Error Comparison on Synthetic Datasets
@@
Figure 4. Comparison on synthetic datasets.
Figure 5. Condition TraCS-D outperforms TraCS-C.
Figure 6. TraCS-D vs $k$-$R$ + uniform sampling.
Figure 4 presents the error comparison between TraCS and the strawman method. We can observe the error advantage of TraCS over the truncated Laplace mechanism (T-Laplace) across all $\varepsilon$ values, and over the strawman method as the privacy parameter $\varepsilon$ increases. Specifically, in Figure 4(a), TraCS-C exhibits smaller AE when $\varepsilon\gtrsim 3$, and TraCS-D exhibits smaller AE when $\varepsilon\gtrsim 5$. When the location space expands to $0,2)\times[0,10)$, i.e. Figure 4(b), the advantage of TraCS becomes more significant. Among TraCS, TraCS-C consistently has a smaller AE than TraCS-D. This is because TraCS-D often deals with larger distance spaces $\mathcal{D}_{r(\varphi)}$, which also magnifies the error of direction perturbation. Statistically, the mean AE of TraCS-D is $91.1\%$ of the strawman method in Figure 4(a) and $86.6\%$ in Figure 4(b) across all the $\varepsilon$ values. For TraCS-C the corresponding ratios are $75.5\%$ in Figure 4(a) and $64.0\%$ in Figure 4(b).
@@
Figure 7. Comparison on real-world datasets.
4.1.3. Error Comparison on Real-world Datasets
Figure 7 shows the error comparison on the TKY and CHI areas. The trends largely mirror those observed on the synthetic datasets. T-Laplace incurs substantially larger AE than TraCS and the strawman method on the TKY dataset. Across all $\varepsilon$ values, the mean AE of TraCS-D is $96.8\%$ of the strawman method on TKY and $94.5\%$ on CHI, for TraCS-C the corresponding ratios are $89.7\%$ and $61.2\%$. The performance gap is smaller on TKY than on CHI, because TKY’s location space is more compact, which lowers the error values for all perturbation methods.
4.1.4. Choose TraCS-D or TraCS-C
@@
4.2. Results on Discrete Space
Our approach can be applied to discrete spaces by rounding each perturbed location to its nearest discrete point. Importantly, the privacy guarantee is established in the underlying continuous domain and therefore still holds after rounding, regardless of the chosen discretization. In the following, we evaluate the performance of TraCS on discrete spaces and compare it with existing methods.
4.2.1. Setup
@@
  • NGram (DBLP:journals/pvldb/CunninghamCFS21): It is often impractical to acquire reachability knowledge in practice (DBLP:journals/pvldb/Zhang [^0] H23). We instead consider a strong reachability constraint, i.e. from each location, the next location can only be within a distance of $0.75\cdot\hat{a}$, [^10] While the original NGram employs much stricter time-reachability constraints (e.g. the next location must be within 0.10.1h $\cdot$ 44km/h), this constraint is also known to the adversary, which undermines the privacy guarantee. TraCS are LDP mechanisms guaranteeing privacy for the entire domain and do not incorporate such constraints. where $\hat{a}$ is the maximal length of the location space. This constraint significantly reduces the location space for perturbation, especially for the locations near the boundary.
  • L-SRR (DBLP:conf/ccs/WangH0QH22): We use group number $g=3$, the same as the empirical optimal value in their paper and code. Locations in the first group are within $0.3\cdot\hat{a}$ distance from the sensitive location, where $\hat{a}$ is the maximal length of the location space. The second group is within $0.6\cdot\hat{a}$ distance from the sensitive location, and the remaining locations are in the third group.
  • ATP (DBLP:journals/pvldb/Zhang [^0] H23): We set the $k$-$R$ parameter $k=6$ in ATP. The privacy parameter $\varepsilon$ in original ATP is for the entire trajectory, we distribute it to each location’s perturbation according to their algorithms.
@@
  • Synthetic datasets: We uniformly discretize $0,1)\times[0,1)$ into $m$ discrete points and treat the $m$ points as the discrete location space. Specifically, we set $m=10\times 10$ and $m=60\times 60$ for comparison. Along with them, we generate $100$ random trajectories with $100$ locations each as synthetic datasets, i.e. each location is randomly sampled from the $m$ discrete points and connected sequentially.
  • TKY and CHI (visualized in Figure 18): We treat every location in the TKY and CHI datasets as a discrete point, which together constitute the discrete location space. TKY contains $7,798$ discrete locations; CHI contains $1,000$ discrete locations. We use the first $100$ trajectories for evaluation.
For the TKY and CHI datasets, we perform the perturbation $5$ times for each trajectory and compute the average AE, which is the same setup as in ATP (DBLP:journals/pvldb/Zhang [^0] H23).
4.2.2. Error Comparison on Synthetic Datasets
@@
Figure 8. Comparison on synthetic datasets. $m$ is the number of discrete points in the location space.
@@
4.2.3. Error Comparison on Real-world Datasets
@@
Reasons for the results. (i) ATP performs exceptionally well when $\varepsilon$ is small, particularly on the TKY dataset. This is because this dataset contains many “condensed” locations, i.e. no other locations near the trajectory. In such cases, if the bi-direction perturbation of ATP is accurate, the perturbed location is error-free. (ii) The AE of ATP does not decrease significantly as $\varepsilon$ increases, which is consistent with their paper. NGram shows a similar trend, especially on the TKY dataset. This is due to the low performance of the Exponential mechanism for large location spaces, particularly when the locations are concentrated in a small area. In such cases, the score function struggles to differentiate the scores among the locations. L-SRR exhibits a similar trend. (iii) TraCS exhibits larger AE than them when $\varepsilon$ is small because it focuses on the rectangular area formed by the longitude and latitude, instead of a set of discrete locations in the cities (visualized in Figure 18). Therefore, the perturbed locations may be far from the city center (where benefits discrete mechanisms) when $\varepsilon$ is small.
@@
Figure 9. Comparison on real-world datasets. TraCS operates on rectangular areas encompassing the city, which may result in larger AE when $\varepsilon$ is small.
Table 3. Time cost comparison (in milliseconds).
[…]
4.2.4. Time Cost Comparison